A butcher has some hamburger that is 21% fat and some that is 15% fat. How many pounds of each should be mixed to make 78 lb of hamburger that is 17% fat?

21% fat _?_ lb
15% fat _?_ lb

Answer by lwsshak3(11628)

A butcher has some hamburger that is 21% fat and some that is 15% fat. How many pounds of each should be mixed to make 78 lb of hamburger that is 17% fat?
let x= pounds of 21% hamburger required
78-x =pounds of 15% hamburger required
.21x+.15(78-x)=.17(78)
multiply both sides by 10
21x+15(78-x)=17(78)
21x+1170-15x=1326
6x=1326-1170
6x=156
x=156/6=26 lbs
78-x=78-26=52 lbs
ans: pounds of 21% hamburger required = 26
pounds of 15% hamburger required = 52